# Latent Schrödinger Bridge Diffusion Model for Generative Learning

### v1

@@ Appendix B Pre-training
$|\ell_{\boldsymbol{E},\boldsymbol{D}}(\mathbf{y})-\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y})|$
$\leq$ $(B_{\boldsymbol{D}}+K_{\boldsymbol{D}}+2\sqrt{d})\left(\|(\boldsymbol{D}\circ\boldsymbol{E})(\mathbf{y})-(\boldsymbol{D}_{\delta}\circ\boldsymbol{E})(\mathbf{y})\|+\|(\boldsymbol{D}_{\delta}\circ\boldsymbol{E})(\mathbf{y})-(\boldsymbol{D}_{\delta}\circ\boldsymbol{E}_{\delta})(\mathbf{y})\|\right)$
$\leq$ $(B_{\boldsymbol{D}}+K_{\boldsymbol{D}}+2\sqrt{d})(\delta+\sqrt{d}\gamma_{\boldsymbol{D}}\delta)$
$=$ $(B_{\boldsymbol{D}}+K_{\boldsymbol{D}}+2\sqrt{d})(1+\sqrt{d}\gamma_{\boldsymbol{D}})\delta.$
For simplicity, we denote $\mathcal{N}_{\mathcal{D}}:=\mathcal{N}(\delta,\mathcal{D},\|\cdot\|_{L^{\infty}([0,1]^{d^{*}})})$, $\mathcal{N}_{\mathcal{E}}:=\mathcal{N}(\delta,\mathcal{E},\|\cdot\|_{L^{\infty}([0,1]^{d})})$ and $C(\delta):=(B_{\boldsymbol{D}}+K_{\boldsymbol{D}}+2\sqrt{d})(1+\sqrt{d}\gamma_{\boldsymbol{D}})\delta$, then
@@
$\leq\max_{\boldsymbol{E}_{\delta}\in\mathcal{E}_{\delta},\boldsymbol{D}_{\delta}\in\mathcal{D}_{\delta}}\left|\frac{1}{\mathcal{M}}\sum_{i=1}^{\mathcal{M}}\left(\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})-\mathbb{E}_{\mathcal{Y}}\left[\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})\right]\right)\right|+2C_{\delta}.$
@@
$\sup_{\boldsymbol{E}\in\mathcal{E},\boldsymbol{D}\in\mathcal{D}}|\hat{\mathcal{H}}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E},\boldsymbol{D})|\leq\max_{\boldsymbol{E}_{\delta}\in\mathcal{E}_{\delta},\boldsymbol{D}_{\delta}\in\mathcal{D}_{\delta}}\left|\frac{1}{\mathcal{M}}\sum_{i=1}^{\mathcal{M}}\left(\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})-\mathbb{E}_{\mathcal{Y}}\left[\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})\right]\right)\right|+2C_{\delta}.$
Thus, for any $t>2C_{\delta}$, we have
$\mathbb{P}_{\mathcal{Y}}\left(\sup_{\boldsymbol{E}\in\mathcal{E},\boldsymbol{D}\in\mathcal{D}}|\hat{\mathcal{H}}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E},\boldsymbol{D})|>t\right)$
$\leq$ $\mathbb{P}_{\mathcal{Y}}\left(\max_{\boldsymbol{E}_{\delta}\in\mathcal{E}_{\delta},\boldsymbol{D}_{\delta}\in\mathcal{D}_{\delta}}\left|\frac{1}{\mathcal{M}}\sum_{i=1}^{\mathcal{M}}\left(\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})-\mathbb{E}_{\mathcal{Y}}\left[\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})\right]\right)\right|>t-2C_{\delta}\right)$
$\leq$ $\sum_{\boldsymbol{D}_{\delta}\in\mathcal{D}_{\delta}}\sum_{\boldsymbol{E}_{\delta}\in\mathcal{E}_{\delta}}\mathbb{P}_{\mathcal{Y}}\left(\left|\frac{1}{\mathcal{M}}\sum_{i=1}^{\mathcal{M}}\left(\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})-\mathbb{E}_{\mathcal{Y}}\left[\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})\right]\right)\right|>t-2C_{\delta}\right).$
@@
$\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y})\leq 2\left(\|(\boldsymbol{D}_{\delta}\circ\boldsymbol{E}_{\delta})(\mathbf{y})\|^{2}+\|\mathbf{y}\|^{2}\right)\leq 2(B_{\boldsymbol{D}}^{2}+d),$
@@
$\mathbb{P}_{\mathcal{Y}}\left(\left|\frac{1}{\mathcal{M}}\sum_{i=1}^{\mathcal{M}}\left(\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})-\mathbb{E}_{\mathcal{Y}}\left[\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})\right]\right)\right|>t-2C_{\delta}\right)\leq 2\exp\left(-\frac{\mathcal{M}(t-2C_{\delta})^{2}}{2(B_{\boldsymbol{D}}^{2}+d)^{2}}\right).$
Therefore, for any $t>2C_{\delta}$, we have
$\mathbb{P}_{\mathcal{Y}}\left(\sup_{\boldsymbol{E}\in\mathcal{E},\boldsymbol{D}\in\mathcal{D}}|\hat{\mathcal{H}}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E},\boldsymbol{D})|>t\right)\\ \leq 2\mathcal{N}_{\mathcal{E}}\mathcal{N}_{\mathcal{D}}\exp\left(-\frac{\mathcal{M}(t-2C_{\delta})^{2}}{2(B_{\boldsymbol{D}}^{2}+d)^{2}}\right).$
@@
$\mathbb{E}_{\mathcal{Y}}\left(\sup_{\boldsymbol{E}\in\mathcal{E},\boldsymbol{D}\in\mathcal{D}}|\hat{\mathcal{H}}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E},\boldsymbol{D})|\right)$
$=$ $\int_{0}^{+\infty}\mathbb{P}_{\mathcal{Y}}\left(\sup_{\boldsymbol{D}\in\mathcal{D},\boldsymbol{E}\in\mathcal{E}}|\hat{\mathcal{H}}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E},\boldsymbol{D})|>t\right)\mathrm{d}t$
$\leq a+2C_{\delta}+2\mathcal{N}_{\mathcal{D}}\mathcal{N}_{\mathcal{E}}\int_{a}^{+\infty}\exp\left(-\frac{\mathcal{M}t^{2}}{2(B_{\boldsymbol{D}}^{2}+d)^{2}}\right)\mathrm{d}t$
$\leq a+2C_{\delta}+2\mathcal{N}_{\mathcal{D}}\mathcal{N}_{\mathcal{E}}\cdot\frac{\sqrt{\pi}}{2}\exp\left(-\frac{\mathcal{M}a^{2}}{2(B_{\boldsymbol{D}}^{2}+d)^{2}}\right)\frac{2(B_{\boldsymbol{D}}^{2}+d)}{\sqrt{2\mathcal{M}}}.$
Taking $a=2(B_{\boldsymbol{D}}^{2}+d)\sqrt{\frac{\log\mathcal{N}_{\mathcal{D}}\mathcal{N}_{\mathcal{E}}}{2\mathcal{M}}}$ and $\delta=\frac{1}{\mathcal{M}}$, we obtain
$\mathbb{E}_{\mathcal{Y}}\left(\sup_{\boldsymbol{E}\in\mathcal{E},\boldsymbol{D}\in\mathcal{D}}|\hat{\mathcal{H}}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E},\boldsymbol{D})|\right)$
$\leq$ $2(B_{\boldsymbol{D}}^{2}+d)\frac{\sqrt{\log\mathcal{N}_{\mathcal{D}}\mathcal{N}_{\mathcal{E}}}+\sqrt{\pi}}{\sqrt{2\mathcal{M}}}+\frac{2(B_{\boldsymbol{D}}+K_{\boldsymbol{D}}+2\sqrt{d})(1+\sqrt{d}\gamma_{\boldsymbol{D}})}{\mathcal{M}}.$
Substituting the values of $\mathcal{N}_{\mathcal{E}}$ and $\mathcal{N}_{\mathcal{D}}$, we obtain
$\mathbb{E}_{\mathcal{Y}}\left(\sup_{\boldsymbol{E}\in\mathcal{E},\boldsymbol{D}\in\mathcal{D}}|\hat{\mathcal{H}}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E},\boldsymbol{D})|\right)=\widetilde{\mathcal{O}}\left(\frac{\epsilon^{-d/2}}{\sqrt{\mathcal{M}}}\right).$
@@
$\mathcal{H}(\hat{\boldsymbol{E}},\hat{\boldsymbol{D}})-\mathcal{H}(\boldsymbol{E}^{*},\boldsymbol{D}^{*})=\widetilde{\mathcal{O}}\left(\epsilon+\frac{\epsilon^{-d/2}}{\sqrt{\mathcal{M}}}\right).$
@@
$\mathcal{H}(\hat{\boldsymbol{E}},\hat{\boldsymbol{D}})-\mathcal{H}(\boldsymbol{E}^{*},\boldsymbol{D}^{*})=\widetilde{\mathcal{O}}\left(\mathcal{M}^{-\frac{1}{d+2}}\right),$

### v2

@@ Appendix A Pre-training
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ |\ell_{\boldsymbol{E},\boldsymbol{D}}(\mathbf{y})-\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y})|$
$\leq(2K_{\boldsymbol{D}}+2\sqrt{d})\left(\|(\boldsymbol{D}\circ\boldsymbol{E})(\mathbf{y})-(\boldsymbol{D}_{\delta}\circ\boldsymbol{E})(\mathbf{y})\|+\|(\boldsymbol{D}_{\delta}\circ\boldsymbol{E})(\mathbf{y})-(\boldsymbol{D}_{\delta}\circ\boldsymbol{E}_{\delta})(\mathbf{y})\|\right)$
$\leq 2(K_{\boldsymbol{D}}+\sqrt{d})(\delta+\sqrt{d}\gamma_{\boldsymbol{D}}\delta)$
$=2(K_{\boldsymbol{D}}+\sqrt{d})(1+\sqrt{d}\gamma_{\boldsymbol{D}})\delta.$
For simplicity, we denote $\mathcal{N}_{\mathcal{D}_{0},\delta}:=\mathcal{N}(\delta,\mathcal{D}_{0},\|\cdot\|_{L^{\infty}([-1,1]^{d^{*}})})$, $\mathcal{N}_{\mathcal{E}_{0},\delta}:=\mathcal{N}(\delta,\mathcal{E}_{0},\|\cdot\|_{L^{\infty}([-1,1]^{d})})$ and $C(\delta):=2(K_{\boldsymbol{D}}+\sqrt{d})(1+\sqrt{d}\gamma_{\boldsymbol{D}})\delta$, then
@@
$\leq\max_{\boldsymbol{E}_{\delta}\in\mathcal{E}_{\delta},\boldsymbol{D}_{\delta}\in\mathcal{D}_{\delta}}\left|\frac{1}{\mathcal{M}}\sum_{i=1}^{\mathcal{M}}\left(\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})-\mathbb{E}_{\mathcal{Y}}\left[\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})\right]\right)\right|+2C({\delta}).$
@@
$\sup_{\boldsymbol{E}\in\mathcal{E}_{0},\boldsymbol{D}\in\mathcal{D}_{0}}|\widehat{\mathcal{H}}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E},\boldsymbol{D})|\leq\max_{\boldsymbol{E}_{\delta}\in\mathcal{E}_{\delta},\boldsymbol{D}_{\delta}\in\mathcal{D}_{\delta}}\left|\frac{1}{\mathcal{M}}\sum_{i=1}^{\mathcal{M}}\left(\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})-\mathbb{E}_{\mathcal{Y}}\left[\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})\right]\right)\right|+2C({\delta}).$
Thus, for any $t>2C({\delta})$, we have
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{P}_{\mathcal{Y}}\left(\sup_{\boldsymbol{E}\in\mathcal{E}_{0},\boldsymbol{D}\in\mathcal{D}_{0}}|\widehat{\mathcal{H}}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E},\boldsymbol{D})|>t\right)$
$\leq\mathbb{P}_{\mathcal{Y}}\left(\max_{\boldsymbol{E}_{\delta}\in\mathcal{E}_{\delta},\boldsymbol{D}_{\delta}\in\mathcal{D}_{\delta}}\left|\frac{1}{\mathcal{M}}\sum_{i=1}^{\mathcal{M}}\left(\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})-\mathbb{E}_{\mathcal{Y}}\left[\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})\right]\right)\right|>t-2C({\delta})\right)$
$\leq\sum_{\boldsymbol{D}_{\delta}\in\mathcal{D}_{\delta}}\sum_{\boldsymbol{E}_{\delta}\in\mathcal{E}_{\delta}}\mathbb{P}_{\mathcal{Y}}\left(\left|\frac{1}{\mathcal{M}}\sum_{i=1}^{\mathcal{M}}\left(\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})-\mathbb{E}_{\mathcal{Y}}\left[\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})\right]\right)\right|>t-2C({\delta})\right).$
@@
$\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y})\leq 2\left(\|(\boldsymbol{D}_{\delta}\circ\boldsymbol{E}_{\delta})(\mathbf{y})\|^{2}+\|\mathbf{y}\|^{2}\right)\leq 2(K_{\boldsymbol{D}}^{2}+d),$
@@
$\mathbb{P}_{\mathcal{Y}}\left(\left|\frac{1}{\mathcal{M}}\sum_{i=1}^{\mathcal{M}}\left(\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})-\mathbb{E}_{\mathcal{Y}}\left[\ell_{\boldsymbol{E}_{\delta},\boldsymbol{D}_{\delta}}(\mathbf{y}_{i})\right]\right)\right|>t-2C_{\delta}\right)\leq 2\exp\left(-\frac{\mathcal{M}(t-2C({\delta}))^{2}}{2(K_{\boldsymbol{D}}^{2}+d)^{2}}\right).$
Therefore, for any $t>2C({\delta})$, we have
$\mathbb{P}_{\mathcal{Y}}\left(\sup_{\boldsymbol{E}\in\mathcal{E}_{0},\boldsymbol{D}\in\mathcal{D}_{0}}|\widehat{\mathcal{H}}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E},\boldsymbol{D})|>t\right)\\ \leq 2\mathcal{N}_{\mathcal{E}_{0},\delta}\mathcal{N}_{\mathcal{D}_{0},\delta}\exp\left(-\frac{\mathcal{M}(t-2C({\delta}))^{2}}{2(K_{\boldsymbol{D}}^{2}+d)^{2}}\right).$
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{E}_{\mathcal{Y}}\left(\sup_{\boldsymbol{E}\in\mathcal{E}_{0},\boldsymbol{D}\in\mathcal{D}_{0}}|\widehat{\mathcal{H}}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E},\boldsymbol{D})|\right)$
$=\int_{0}^{+\infty}\mathbb{P}_{\mathcal{Y}}\left(\sup_{\boldsymbol{D}\in\mathcal{D}_{0},\boldsymbol{E}\in\mathcal{E}_{0}}|\widehat{\mathcal{H}}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E},\boldsymbol{D})|>t\right)\mathrm{d}t$
$\leq a+2C({\delta})+2\mathcal{N}_{\mathcal{D}_{0},\delta}\mathcal{N}_{\mathcal{E}_{0},\delta}\int_{a}^{+\infty}\exp\left(-\frac{\mathcal{M}t^{2}}{2(K_{\boldsymbol{D}}^{2}+d)^{2}}\right)\mathrm{d}t$
$\leq a+2C({\delta})+2\mathcal{N}_{\mathcal{D}_{0},\delta}\mathcal{N}_{\mathcal{E}_{0},\delta}\cdot\frac{\sqrt{\pi}}{2}\exp\left(-\frac{\mathcal{M}a^{2}}{2(K_{\boldsymbol{D}}^{2}+d)^{2}}\right)\frac{2(K_{\boldsymbol{D}}^{2}+d)}{\sqrt{2\mathcal{M}}}.$
Taking $a=2(K_{\boldsymbol{D}}^{2}+d)\sqrt{\frac{\log\mathcal{N}_{\mathcal{D}_{0},\delta}\mathcal{N}_{\mathcal{E}_{0},\delta}}{2\mathcal{M}}}$ and $\delta=\frac{1}{\mathcal{M}}$, we obtain
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{E}_{\mathcal{Y}}\left(\sup_{\boldsymbol{E}\in\mathcal{E}_{0},\boldsymbol{D}\in\mathcal{D}_{0}}|\widehat{\mathcal{H}}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E},\boldsymbol{D})|\right)$
$\leq 2(K_{\boldsymbol{D}}^{2}+d)\frac{\sqrt{\log\mathcal{N}_{\mathcal{D}_{0},\delta}\mathcal{N}_{\mathcal{E}_{0},\delta}}+\sqrt{\pi}}{\sqrt{2\mathcal{M}}}+\frac{4(K_{\boldsymbol{D}}+\sqrt{d})(1+\sqrt{d}\gamma_{\boldsymbol{D}})}{\mathcal{M}}.$
Substituting the values of $\mathcal{N}_{\mathcal{E}_{0},\delta}$ and $\mathcal{N}_{\mathcal{D}_{0},\delta}$, we obtain
$\mathbb{E}_{\mathcal{Y}}\left(\sup_{\boldsymbol{E}\in\mathcal{E}_{0},\boldsymbol{D}\in\mathcal{D}_{0}}|\widehat{\mathcal{H}}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E},\boldsymbol{D})|\right)=\widetilde{\mathcal{O}}\left(\frac{\epsilon^{-d/2}}{\sqrt{\mathcal{M}}}\right).$
@@
$\mathcal{H}(\widehat{\boldsymbol{E}},\widehat{\boldsymbol{D}})-\mathcal{H}(\boldsymbol{E}^{*},\boldsymbol{D}^{*})=\widetilde{\mathcal{O}}\left(\epsilon+\frac{\epsilon^{-d/2}}{\sqrt{\mathcal{M}}}\right).$
@@
$\mathcal{H}(\widehat{\boldsymbol{E}},\widehat{\boldsymbol{D}})-\mathcal{H}(\boldsymbol{E}^{*},\boldsymbol{D}^{*})=\widetilde{\mathcal{O}}\left(\mathcal{M}^{-\frac{1}{d+2}}\right),$